A 104.2 mL sample of 1.00 M NaOH is mixed with 52.1 mL of 1.00 M H2SO4 in a large Styrofoam coffee cup; the cup is fitted with a lid through which passes a calibrated thermometer. The temperature of each solution before mixing is 23.05 °C. After adding the NaOH solution to the coffee cup and stirring the mixed solutions with the thermometer, the maximum temperature measured is 32.30 °C. Assume that the density of the mixed solutions is 1.00 g/mL, that the specific heat of the mixed solutions is 4.18 J/(g·°C), and that no heat is lost to the surroundings.

Calculate the enthalpy change per mole of H2SO4 in the reaction
Answer #1

Heat liberated(q) = m*s*DT

m = mass of reaction mixture = 104.2+52.1 = 156.3 g ( density of mixture = 1 g/cc)

s = specific heat of mixture = 4.18 j/g.c

DT = 32.3-23.05 = 9.25

q = 156.3*4.18*9.25 = 6043.34 joule

q = 6.0 kj

n = No of mole of H2SO4 reacted = 52.1*1/1000 = 0.0521 mole

DHrxn = - q/n

      = - 6.0/0.0521

       = -115.2 kj/mol